OpenStudy (anonymous):

Write the slope-intercept form of the equation for the line. (-4,-2), (4, 3) I think it's y = -5/8x + 1/2 But i'm not sure. I'm probably wrong, i'm really bad when it comes to slopes.

Mehek (mehek14):

let me check

Mehek (mehek14):

did you use \(\Huge \dfrac{y_2 - y_1}{x_2 - x_1}\)?

OpenStudy (anonymous):

Yeah. But I could still have done the math wrong. Happens lol.

Mehek (mehek14):

\(\dfrac{3 - (-2)}{4 - (-4)}\)

Mehek (mehek14):

it is \(\dfrac{5}{8}\) no negative sign

Mehek (mehek14):

now i check the y-int

OpenStudy (anonymous):

Thannnkkkk you c:

Mehek (mehek14):

3 = 5/8 x 4 + b 3 = 2.5 + b 0.5 = b

Mehek (mehek14):

so yea you were correct just no negative sign

OpenStudy (anonymous):

I appreciate the help c: So much. Thank you.

Mehek (mehek14):

ur welcome :) anytime